Assume that you have a committee of 13 members, made up of 6 men and 7 women.
In how many ways can the 3 tasks be assigned so that both men and women are given assignments?

Answer by stanbon(75887)

Assume that you have a committee of 13 members, made up of 6 men and 7 women.
In how many ways can the 3 tasks be assigned so that both men and women are given assignments?
-------
# of ways to select 3 women: 7C3
# of ways to select 3 men: 6C3
-----
Total # of ways to select 3: 13C3
----
# of ways to select 3 so both men and
and women are given assignments:
13C3 - 7C3 - 6C3 = 286 - 35 - 20 = 286-55 = 281 ways
==================
Cheers,
Stan H.
==================